# Morpheus: Benchmarking Physical Reasoning of Video Generative Models with Real Physical Experiments

### v1

@@ 3.3 Trajectory extraction
While generated videos could be directly evaluated in terms of 3D consistency [19] or other pixel-level generation properties [5], such evaluations are limited to visual and geometric realism of the generated videos. Instead, we are interested in how well these videos conform to physical laws. This means that we need to extract the relevant physical state variables, such as positions of objects, velocities, accelerations, masses, and so on. Thus, it is essential to transform the generated videos into perfect state variables of the depicted objects and their trajectories, which can then be further analyzed.
A supervised object-tracking model would work well for specific video domains. However, for unseen generated videos, we need more general methods for reliable processing. Thus, we propose leveraging a vision foundation model for zero-shot object segmentation and tracking. In particular, we use Segment Anything 2 (SAM2) [34], a promptable segmentation and tracking model for images and videos that provides 2D masks. We label the first frames of the videos in our dataset using positive and negative clicks as spatial prompts and verify the segmentation before tracking the object throughout the video. Finally, we extract center-of-mass 2D coordinates by averaging object mask centers. In addition, to check if objects have consistent depth (necessary for using only 2D coordinates for our physical scores, see App. C.2), we estimate relative object depth using Depth Anything V2 [35] to predict depth values for the corresponding masks from SAM2 (see Fig. 3 for illustration).
For velocity, acceleration and angular velocity, we employ the central difference method  [36]. To further reduce the noise, generated by the imperfections in the tracking pipeline, we follow with a series of smoothing operations, such as learning a linear regression with a sliding window and applying the Savitzky-Golay smoothing. The details can be found in the App. B.
Since the generated videos may contain artifacts that prohibit further analysis, e.g. objects’ permanence, jittering, or absence of movement, we discard such samples and keep track of the discard rate for each model.
Figure 5: Evaluation of trajectories extracted from real and VGMs videos using Dynamical score. For each of the trajectory, real-world or generated videos, we fit a PINN with the corresponding equation of motion for the particular physical law as an extra loss term.
@@ 4 Physics-informed evaluation metrics
As a first metric, we compute the discard rate, which reflects the proportion of model-generated samples that must be discarded to ensure reliable trajectory extraction needed for Physical Invariances and Dynamical scores. The discard filtering is automatic and consists of three criteria: First, we discard generated videos where objects lack sufficient permanence throughout the video. Second, we discard generated videos which do not have a consistent number of objects. Finally, we discard generated videos if there is little motion detected, as such videos are not suitable for physical analysis. The overall discard rate represents the proportion of generated videos that fail at least one of these criteria. In addition, we verify that none of the real-world extracted trajectories are discarded, showing that our discard criteria are effective in distinguishing physical from non-physical videos. We provide further details on our filtering methodology in App. C.1.
Beyond pixel-by-pixel evaluation
For the videos that pass the filtering, we employ two metrics: the Dynamical score, which measures the overall adherence to the equation of motion that governs the system, and the Physical Invariance score, which quantifies the invariance of conserved physical quantities, such as energy or angular momentum (Table A2).
As discussed above, pixel-by-pixel evaluation is inherently problematic when perfect control over the initial conditions is not achievable. Even with ideal control, the chaotic nature of certain physical phenomena can cause minor differences in initial conditions to produce vastly divergent outcomes. Consequently, straightforward time-based trajectory matching becomes an unreliable measure of performance. For our benchmark, we instead rely on metrics that do not depend on direct time-aligned comparisons. These metrics are designed to evaluate physical consistency and adherence to fundamental laws independently of precise time synchronization, providing a more robust and generalizable standard of evaluation. In our approach, the obtained real-world trajectories act as reference points to establish the upper bounds of performance and validate our methods.
@@ 4.1 Dynamical score
To calculate the Dynamical score, we use physics-informed neural networks (PINNs)[14], which directly incorporate physical laws as a prior. This setting allows us to learn the physical trajectory that fits the data the most, independent of the initial conditions. Fig. 5 illustrates our approach. A PINN is a neural network that receives a timestep $i$ of the trajectory as input and outputs the trajectory coordinates $\hat{T_{i}}$, velocity $\hat{\dot{T_{i}}}$, and acceleration $\hat{\ddot{T_{i}}}$. The model is typically trained with a loss function, comprising two components $L_{\text{data}}$ and $L_{\text{physics}}$: $L_{\text{total}}=L_{\text{data}}+\lambda L_{\text{physics}}$, where the $L_{\text{data}}$ is responsible for fitting the model to the datapoints, $L_{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{T}_{i}-T_{i}\|^{2}$, and $L_{\text{physics}}$ enforces following the physical law. For each experiment, we explicitly implement the equation of motion in the form of an ordinary differential equation as PINN loss functions. E.g., for the falling ball, the equations of motion are: $\dot{x}=0;\quad\ddot{y}+g=0$, where $y$ is the vertical position, $\ddot{y}$ is the acceleration and $g$ is the gravitational constant. The $L_{\text{physics}}$ is calculated as: $L_{\text{physics}}=\frac{1}{M}\sum_{j=1}^{M}\left\|\hat{\ddot{y}}_{j}+g\right\|^{2}+\left\|\hat{\dot{x}}\right\|^{2}$, where $\hat{\ddot{y}}_{j}$ is the predicted acceleration derived from the PINN at the $j$-th time step.
@@
In our context, we use PINNs to assess the physical plausibility of trajectories from generated videos by computing the normalized mean square error (NMSE) of the model-learned trajectory derived from real and generated videos. To normalize it into the range of $(0,1)$, we inverse the error with $1$ indicating a maximal Dynamical score and 00 indicating worse than constant function fit of the corresponding PINN. Dynamical score shows the difference between theoretical prediction from the ODEs and trajectory data. A higher Dynamical score implies higher physical plausibility. For more details, please see App. C.3.
@@ 4.2 Physical Invariance score
To calculate a more fine-grained Physical Invariance score, we accompany each of our experiments with a list of physical invariances, i.e. values that we can derive from trajectories that stay constant in time. For the physical model to work, we make a series of reasonable assumptions about the setting and test them on the real-world trajectories. As invariances vary for different experiments, we present here one case study for the falling ball experiments, while describing all the other settings in App. C.4. We list in Table A2 all theoretically conserved values per experimental scenario.
@@
$\Rightarrow$ Energy-to-mass ratio. Assuming that the mass of the ball is constant, we derive the following invariant:$\frac{E}{m}=\frac{1}{2}(v_{x}^{2}+v_{y}^{2})+gy=\text{const}$, which we can estimate with the data from our trajectory.
$\Rightarrow$ Acceleration. As no external forces are acting on the ball except for gravity, which is uniform in space and time and is directed downwards, the acceleration of the ball is also constant: $a_{y}=g=\text{const}$.
Figure 6: Physical Invariance vs. Dynamical scores for all the experiments. Different colors correspond to different models, whereas the hue of the color differentiates between base and enhanced prompts. Scores from trajectories extracted from real-world videos are consistently close to maximum scores, while deviations are explained by trajectory estimation errors and using only 2D trajectories.
Figure 7: Average discard rates across all physical experiments.
$\Rightarrow$ Horizontal momentum-to-mass ratio. As with acceleration, the horizontal momentum, $p_{x}=mV_{x}$, is also preserved given no external forces. This implies that the horizontal velocity is conserved: $v_{x}=\text{const}$
@@
To convert the invariant into an actual score, like the Energy score, we calculate the standard deviation of the invariant time series and normalize it into the range of $(0,1)$, with 1 indicating a perfect Physical Invariance score.
As invariants must be by nature constant, a high standard deviation of these invariants (and thus a lower physical invariance score), indicates poor modeling of the respective physical invariants. In addition, for discarded trajectories we assign minimal Physical Invariance score equal to 00. A detailed score calculation procedure is described in the App. C.5. For the derivations of each invariant we used, please refer to the App. C.4.
Figure 8: Average Physical Invariance vs. Dynamical scores over the experiments. Different colors correspond to different models, whereas the hue of the color differentiates between base and enhanced prompts.
@@ 5 Analysis
In this section, we analyze the results of the experiments. In tables A5 to A9, as well as Fig. 6 and Fig. 8, we note plain for simple text prompts, and enhanced for upgraded textual description (see Sec.3.2). We consider the multi-frame prompting scenario separately.
@@
Enhanced prompts typically improve performance metrics compared to plain prompts, although this trend varies depending on the specific model. Enhanced prompting leads to higher physical invariance and dynamical scores and lower discard rates in many cases (e.g., COSMOS and CogVideoX), yet occasionally decreases performance in models such as LTX and PyramidalFlow, indicating that prompt enhancement effectiveness is context-dependent.
@@
Multi-frame prompting generally outperforms single-frame prompting, achieving lower discard rates and higher dynamical and physical invariance scores, thereby demonstrating the advantage of increased temporal context. The prompting with first and last frames, used exclusively by CogVideoX as an interpolation regime, performs notably well in specific experiments (e.g., holonomic pendulum), suggesting a promising direction for improving temporal coherence and physical realism, though limiting the ability to generate from scratch.
Among the evaluated models, CogVideoX typically demonstrates superior performance, especially in multi-frame configurations with enhanced prompts. LTX, while occasionally excelling in specific scenarios (such as single-frame enhanced prompts for the falling ball), exhibits inconsistent performance, with high variability and notably high discard rates in certain tasks. COSMOS and PyramidalFlow show intermediate performance, with COSMOS performing strongly in multi-frame scenarios and PyramidalFlow achieving moderate results.
@@
The variability of discard rates across setups reflects the reliability of different models in generating physically plausible videos. Discard rates vary significantly, with extremes ranging from as low as $0.0\%$ (COSMOS single-frame, plain prompt, double pendulum) to as high as $92\%$ (LTX single-frame, enhanced prompt, holonomic pendulum). The analysis of the major reasons (see Fig.7) behind high discard rates reveals the absence of motion (i.e. stillness) and the presence of duplicate objects, as well as, to a lesser extent, the disappearance of the object from the video. These persistent shortcomings in the models’ abilities to produce consistent and realistic videos are well-known [21].
Analyzing fine-grained conservation metrics across experiments reveals several interesting trends. In the falling and bouncing ball scenarios, relatively high horizontal momentum conservation scores can primarily be attributed to the absence of significant horizontal motion rather than accurate modeling of dynamics. In contrast, for the projectile task, where horizontal motion is inherently present, horizontal momentum conservation scores align closely with the generally lower scores observed for energy and acceleration conservation, reflecting the genuine complexity of modeling horizontal dynamics. Additionally, multi-frame prompting substantially enhances distance conservation metrics, likely because maintaining spatial relationships between objects across frames is fundamentally simpler than modeling intricate dynamical properties. Moderate improvements in energy and period conservation further support the advantage of multi-frame temporal context. In pendulum scenarios, some periodic behavior emerges in generated videos (see Fig. 9 and Fig. A10), suggesting that models partly capture periodicity; however, adherence to true dynamical behavior remains limited, as demonstrated by, for example, the low-energy conservation scores.
@@
Overall, all generated models exhibit substantial limitations compared to real-world performance (see Fig. 8), underscoring the significant gaps remaining in simulating realistic and physically accurate dynamics. We present additional per experiment and prompt analysis in the App. F.
@@ 6 Limitations
As the first to analyze detailed physics invariances in generative models, we note individual scores can be misleading alone. For example, a levitating object might preserve energy (appearing correct) while violating gravity. Thus, Physical Invariance scores must be combined with Dynamical ones to get a complete and detailed picture of generative model performance.
@@ 7 Conclusion
Our study highlights a fundamental limitation in current video generation models: despite their impressive realism, they fail to consistently adhere to physical laws. To address this gap, we introduced Morpheus, a benchmark designed to assess the physical reasoning capabilities of these models. Through a curated dataset of real-world physics experiments and physics-informed evaluation metrics, we demonstrate that even with advanced prompting techniques, existing models struggle to capture fundamental physical principles. In general, all models perform poorly, with significant violations of physical principles, though multi-frame prompting provides some improvement. This underscores the need for future research in integrating physical constraints into generative models. We open-source our dataset, baselines, and code to foster further advancements in physics-aware video generation at physics-from-video.github.io/morpheus-bench.
Acknowledgments
AZ is funded by the European Union (ERC, EVA, 950086). We thank Prof. Maziyar Jalaal for providing the equipment for physical experiments.
@@ Acknowledgments
CZ came up with the idea of measuring the physical reasoning ability of video generation models. EG, DC, AZ and CZ set up the initial meeting to structure the project into three parts: experiments recording, trajectory tracking and extracting, and scores calculation. DC and AT collected all the real-world videos with the help of DP. AT, TN and DP set up video generation code for all models with the help from AZ. AZ set up object segmentation and tracking and integrated all parts of the benchmark into automated pipeline with the help from TN and AT. AV completed the Dynamical scores evaluation, while CZ designed the Physical Invariant Scores. The paper was written by CZ, DC, AZ and EG with DC and DP contributing to Fig. 1, 3, 4, 5, AV created Fig. 9 and TN contributed to Fig. 2. AZ created Table 1 and made all the experimental results aggregation and visualization in Fig. 6, 7, 8 with the help from MB. The website development was coordinated by AZ with contributions from CZ (leaderboard), DC (videos), MB (overall structure and plots).

### v2

@@ 3.2 Trajectory extraction
While generated videos could be directly evaluated in terms of 3D consistency (Liu et al., 2024a) or other pixel-level generation properties (Agarwal et al., 2025), such evaluations are limited to visual and geometric realism of the generated videos. Instead, we are interested in how well these videos conform to physical laws. This means that we need to extract the relevant physical state variables, such as positions of objects, velocities, accelerations, masses, and so on. Thus, it is essential to transform the generated videos into perfect state variables of the depicted objects and their trajectories, which can then be further analyzed.
Figure 4: Evaluation of trajectories extracted from real and VGMs videos using our Dynamical (upper) and Physical Invariance (lower) scores. For the Dynamical score, trajectories from real-world or generated videos are fitted to a PINN with the corresponding equation of motion for the particular physical law as an extra loss term. For the Physical Invariance score, using the same trajectories, we estimate physical quantities that should be invariant in the systems, such as total energy and oscillation period, and use their variance as a measure of physical plausibility.
As we need to track objects in both real-world and generated videos, Segment Anything 2 (SAM-2) (Ravi et al., 2024) serve as a reliable way to generate 2D masks for any type of objects. We annotate the first frame of our videos in the dataset with positive and negative labels. Given the masks generated by SAM-2’s we extarct the centroid of the object(s) (center-mass) in the video, at each frame of the video. In addition, we employ Depth Anything V2 (Yang et al., 2024a) to verify that objects have consistent depth through the sequence of movement (See physical score pipeline App. D.2). The depth values are calculated using the correspoding mask generated by SAM-2.
For velocity, acceleration and angular velocity, we employ the central difference method  (Swanson & Turkel, 1992). To further reduce the noise, generated by the imperfections in the tracking pipeline, we follow with a series of smoothing operations, such as learning a linear regression with a sliding window and applying the Savitzky-Golay smoothing. The details can be found in the App. C.
@@ 4 Physics-informed evaluation metrics
As a first metric, we compute the discard rate, which reflects the proportion of model-generated samples that must be discarded to ensure reliable trajectory extraction needed for Physical Invariances and Dynamical scores. The discard filtering is automatic and consists of three criteria: First, we discard generated videos where objects lack sufficient permanence throughout the video. Second, we discard generated videos which do not have a consistent number of objects. Finally, we discard generated videos if there is little motion detected, as such videos are not suitable for physical analysis. The overall discard rate represents the proportion of generated videos that fail at least one of these criteria. In addition, we verify that none of the real-world extracted trajectories are discarded, showing that our discard criteria are effective in distinguishing physical from non-physical videos. We provide further details on our filtering methodology in App. D.1. For the videos that pass filtering, we employ two metrics: Dynamical score, which measures adherence to the governing equation of motion, and Physical Invariance score, which quantifies invariance of conserved quantities such as energy or angular momentum (see Table 3 for all invariances).
Physics beyond trajectory matching
Previous benchmarks such as PhysicsIQ (Motamed et al., 2025) and COSMOS (Alhaija et al., 2025) use trajectory matching by comparing generated trajectories with the ground-truth trajectories. As ground-truth trajectories themselves vary due to noise and hidden physical parameters that are not fully observable from visual image/video conditioning (e.g., object mass or friction), PhysicsIQ proposed to compare the obtained trajectory matching score with the variance in the real-world trajectories. However, such variance depends on the studied /recorded variations and can be arbitrarily large in cases where hidden parameters vary significantly.
Given the visual nature of the conditioning in VGMs (i.e., initial image or video), there are often parameters of the physical system that cannot be estimated from the provided visual conditioning. In the image conditioning case, such unobserved parameters exist for almost all conditions/experiments, limiting the trajectory matching metric for state-of-the-art image-based VGM, such as Veo3 (Veo-Team et al., 2024) and WAN2.1 (Wan et al., 2025a), which allow only image-based conditioning. Moreover, even in the video conditioning case there are often hidden parameters not observed in the original conditioning, limiting the applicability of trajectory matching only to simple systems. In contrast, Morpheus scores overcome those limitations by evaluating results of video generation on the level of physical laws instead of a particular trajectory provided by ground-truth videos.
Figure 5: Different types of discarded generated videos: (left) A video showing the disappearance of the orange ball during fall; (middle) A video illustrating generation of multiple objects in pendulum experiment; (right) A video in which the double pendulum does not move.
@@ 4.1 Dynamical score
To calculate the Dynamical score, we use physics-informed neural networks (PINNs) (Cuomo et al., 2022), which directly incorporate physical laws as a prior. This setting allows us to learn the physical trajectory that fits the data the most, independent of the initial conditions. Fig. 4 illustrates our approach. A PINN is a neural network that receives a timestep $i$ of the trajectory as input and outputs the trajectory coordinates $\hat{T_{i}}$, velocity $\hat{\dot{T_{i}}}$, and acceleration $\hat{\ddot{T_{i}}}$. The model is typically trained with a loss function, comprising two components $L_{\text{data}}$ and $L_{\text{physics}}$: $L_{\text{total}}=L_{\text{data}}+\lambda L_{\text{physics}}$, where the $L_{\text{data}}$ is responsible for fitting the model to the datapoints, $L_{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{T}_{i}-T_{i}\|^{2}$, and $L_{\text{physics}}$ enforces following the physical law. For each experiment, we explicitly implement the equation of motion in the form of an ordinary differential equation as PINN loss functions. E.g., for the falling ball, the equations of motion are: $\dot{x}=0;\quad\ddot{y}+g=0$, where $y$ is the vertical position, $\ddot{y}$ is the acceleration and $g$ is the gravitational constant. The $L_{\text{physics}}$ is calculated as: $L_{\text{physics}}=\frac{1}{M}\sum_{j=1}^{M}\left\|\hat{\ddot{y}}_{j}+g\right\|^{2}+\left\|\hat{\dot{x}}\right\|^{2}$, where $\hat{\ddot{y}}_{j}$ is the predicted acceleration derived from the PINN at the $j$-th time step.
@@
We use PINNs to assess the physical plausibility of trajectories from generated videos by computing the normalized mean square error (NMSE) of the model-learned trajectory derived from videos. We normalize by inverting the error, so $1$ marks the best dynamical score and 0 a worse-than-constant PINN fit. A higher Dynamical score implies higher physical plausibility. For more details, please see App. D.3.
@@ 4.2 Physical Invariance score
To calculate a more fine-grained Physical Invariance score, we accompany each of our experiments with a list of physical invariances, i.e. values that we can derive from trajectories that stay constant in time. We make a series of reasonable assumptions about the setting and test them on the real-world trajectories. As invariances vary for different experiments, we present here one case study for the falling ball experiments, while describing all the other settings in App. D.4 and Table 3.
@@
[…]
$\Rightarrow$ Acceleration. As no external forces are acting on the ball except for gravity, which is uniform and is directed downwards, the acceleration of the ball is also constant: $a_{y}=g=\text{const}$.
$\Rightarrow$ Horizontal momentum-to-mass ratio. As with acceleration, the horizontal momentum, $p_{x}=mV_{x}$, is also preserved given no external forces. Thus, the horizontal velocity is conserved: $v_{x}=\text{const}$.
@@
To convert the invariant into an actual score, like the Energy score, we calculate the standard deviation of the invariant time series and normalize it into the range of $(0,1)$, with 1 indicating a perfect Physical Invariance score. As invariants must be by nature constant, a high standard deviation of these invariants (and thus a lower physical invariance score), indicates poor modeling of the respective physical invariants. In addition, for discarded trajectories we assign minimal Physical Invariance score equal to 0. A detailed score calculation procedure is described in the App. D.5. For the derivations of each invariant we used, please refer to the App. D.4.
@@ 5 Analysis
Figure 6: Average discard rates across all physical experiments (lower is better).
In this section, we analyze the results of the experiments. We present aggregated results for each model and conditioning type (single frame, multi-frame, or key frame interpolation) in Fig. 6 and Fig. 7. For each model, we average the results across all experiments. We select the best scores between using the enhanced (upgraded description, see Sec. 3.1) and plain (simple) versions of the textual prompt to represent the best model’s ability.
@@
Enhanced prompts typically improve performance metrics compared to plain prompts (see Fig. 23), although this trend varies depending on the specific model. Enhanced prompting leads to higher physical invariance and dynamical scores and lower discard rates in many cases (e.g., COSMOS-predict2 and CogVideo), yet occasionally decreases performance in models such as COSMOS-predict1, indicating that prompt enhancement effectiveness is context-dependent.
@@
Multi-frame prompting and key frame interpolation generally outperform single-frame prompting, achieving lower discard rates and higher dynamical and physical invariance scores, thereby demonstrating the advantage of increased temporal context. The prompting with first and last frames, performs notably well in specific experiments (e.g., holonomic pendulum with WAN-2.1, see Fig. 24), suggesting a promising direction for improving temporal coherence and physical realism, though limiting the ability to generate from scratch.
Among the evaluated models, WAN-2.1 with key frame interpolation demonstrates superior performance, with an average total score of 0.55. However, since the last frame is provided, such prompting should be considered as an easier interpolation case vs. hard extrapolation for single and multi-frame conditioning. COSMOS-predict2 shows great results, leading in single frame conditioning category with the score of 0.43, and benefiting greatly from multi-frame conditioning both in reduced discard rates ($9\%$) and overall scores ($0.43\rightarrow 0.52$), and excelling in experiments like bouncing and rolling (see Fig. 18). PyramidalFlow, while occasionally excelling in specific scenarios (e.g., spring, $0.65$), exhibits inconsistent performance, with high variability and notably high discard rates. Other models show intermediate results. LTX can be considered the worst-performing model in the list.
Figure 7: Aggregated scores across all physical experiments (higher is better).
@@
The variability of discard rates across setups reflects the reliability of different models in generating physically plausible videos. Between models, average discard rates vary significantly, ranging from as low as $23.0\%$ (WAN-2.1, key frame interpolation) to as high as $47\%$ (PyramidalFlow, single frame). The variance across experiments is also noticeable, as shown in Fig. 19, Fig. 20, and Fig. 21. Some experiments tend to be prone to only certain types of errors, e.g., stillness in sliding experiments. The analysis of the major reasons (see Fig. 6) behind high discard rates reveals the absence of motion (i.e. stillness) and the presence of duplicate objects, as well as, to a lesser extent, the disappearance of the object from the video. These persistent shortcomings in the models’ abilities to produce consistent and realistic videos are well-known (Huang et al., 2024b).
@@
Overall, all generated models exhibit substantial limitations compared to real-world performance (see Fig. 8), underscoring the significant gaps remaining in simulating physically accurate dynamics.
@@ 6 Limitations
Morpheus is restricted to Newtonian physics under controlled settings, which ensures reproducibility but limits coverage of broader physical processes. To simplify evaluation, we assume negligible air resistance and friction; however, this reduces realism and can penalize physically plausible generations when these assumptions are violated. Next, some invariances are inherently easier to satisfy, making purely observational evaluation less comprehensive. In the future, the benchmark could be extended to cover controllable VGMs, where it would be possible to measure how robust the invariance score is under perturbations. Finally, evaluation remains confined to short videos and simple scenes due to the scope of the vision foundational models’ applicability and the static camera requirement.
@@ 7 Conclusion
Our study highlights a fundamental limitation in current video generation models: despite their impressive realism, they fail to consistently adhere to physical laws. To address this gap, we introduced Morpheus, a benchmark designed to assess the physical reasoning capabilities of these models. Through a curated dataset of real-world physics experiments and physics-informed evaluation metrics, we demonstrate that even with advanced prompting techniques, existing models struggle to capture fundamental physical principles. In general, all models perform poorly, with significant violations of physical principles, though multi-frame prompting provides some improvement. This underscores the need for future research in integrating physical constraints into generative models.
@@
CZ came up with the idea of measuring the physical reasoning ability of video generation models. EG, DC, AZ and CZ set up the initial meeting to structure the project into three parts: experiments recording, trajectory tracking and extracting, and scores calculation. DC and AT collected all the real-world videos with the help of DP. AT, TN and DP set up video generation code for all models with the help from AZ. AZ set up object segmentation and tracking and integrated all parts of the benchmark into automated pipeline with the help from TN and AT. AV completed the Dynamical scores evaluation, while CZ designed the Physical Invariant Scores. The paper was written by CZ, DC, AZ and EG with DC and DP contributing to Fig. 1, 3, 4, 5, AV created Fig. 9, CZ created Fig. 8 and TN contributed to Fig. 2. AZ created Table 1 and made all the experimental results aggregation and visualization in Fig. 6, 7 with the help from MB. TN, AV and DC worked on the final version of the appendix. The website development was coordinated by AZ with contributions from CZ (leaderboard), DC (videos), MB (overall structure and plots).